# Fast Training Dataset Attribution via In-Context Learning

### v1

@@ 1 Introduction
Training Data Attribution (TDA) refers to the task of quantifying contributions of different data sources on the outputs of a model (Park et al., 2023; Nguyen et al., 2023). This task is essential for debugging the processes of curating corpora for training and for improving the training of neural networks. Understanding the contribution of data sources allows us to assess the monetary value of proprietary training data, which is crucial for fair compensation and data management (Ghorbani & Zou, 2019; Nohyun et al., 2022).
Existing methods for TDA, primarily fall into two categories: retraining-based methods and influence function-based methods, as detailed in recent surveys (Hammoudeh & Lowd, 2024; Worledge et al., 2024). Retraining approaches such as those by (Feldman & Zhang, 2020; Ghorbani & Zou, 2019) involve retraining the model without the target data source. However, this method is computationally expensive. Influence function approaches (Koh & Liang, 2017; Pruthi et al., 2020; Chen et al., 2021; Park et al., 2023), relax the need for full retraining by requiring only a few gradient calculations with respect to the data. Despite their efficiency, these methods rely on a linear approximation of the neural network around the target data point, which can be inaccurate. Critically, the influence function approaches compute the attribution score for a dataset as a linear function (usually an average or sum) of the attribution scores for each data point in the dataset (Hammoudeh & Lowd, 2024; Park et al., 2023). This approach fails to provide a holistic view of the contributions of an entire dataset to the model’s output. Additionally, both methods require access to the internals of LLMs, which is not feasible for some popular models. A related technique, Machine Unlearning (Ginart et al., 2019; Sekhari et al., 2021) is still expensive for obtaining the contribution scores.
@@
In the experiments, we evaluate four instruction-tuned LLMs: Mistral 7B (Jiang et al., 2023), Bloomz (Le Scao et al., 2023), Microsoft/Phi-3-mini (Abdin et al., 2024) and GPT 4.0 (Achiam et al., 2023) on a set of binary Q&A datasets, boolq (Clark et al., 2019). We demonstrate that the mixture model approach is more robust to the retrieval noise inherent in RAG systems, providing more accurate estimations of data contributions. Additionally, we conduct an evaluation using Trak (Park et al., 2023) to compare our results with those obtained by the Trak method, validating the effectiveness and efficiency of our proposed approaches.
@@ 2 Methodology
In the binary outcome setting, we have tuples in the format of question, context, and outcome: $(q,c,y)$. When we don’t use any context, we denote $c=\emptyset$. Our model also outputs $p(y|q,c)$. Our goal is to quantify the contributions of the training datasets $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ in $p(y|q,c)$. We assume that we have a query set $Q=\{q_{1},\ldots,q_{m}\}$.
We assume that we have $k,k=1,\ldots,K$ relevant datasets about a topic and we want to quantify their contributions in the generation of the output by our LLM.
@@ 2.1 The Non-parametric Approach: The Shapley Context Method (SCM)
The key idea of this approach is that if an LLM uses the information from the $k$th dataset, providing the $k$th dataset as context will not change the output much. Conversely, if adding a dataset as context changes the output significantly, it was likely not used for generation of the output.
@@
where $c_{k}$ is the context provided from the $k$th dataset.
Usually, desired information can be found in multiple datasets (Ghorbani & Zou, 2019). To consider the impact of datasets in presence of other datasets, we define the following residual scores to be used in the Shapley formula (Shapley, 1953):
$s_{S}=\mathrm{sim}(y\,,y|c_{S}).$ (2)
@@
$\phi_{k}=\sum_{S\subseteq\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}}\frac{|S|!(K-|S|-1)!}{K!}\left(s_{S\cup\{k\}}-s_{S}\right).$ (3)
This formula finds the residual increase in the similarity by including $D_{k}$, when we already have included another set $S\subseteq\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}$. Algorithm 1 in Appendix C describes the details of our Shapley Context Method (SCM).
@@ 2.2 The Semi-Parametric Approach: Context Mixture Factorization (CMF)
We propose a model for summarizing the behavior of LLMs. Our model explicitly contains attribution scores and captures the entirety of the datasets used for its training. We use a mixture distribution approach, which defines:
@@
where $\widetilde{p}_{0}$ denotes a general-purpose language model and $\widetilde{p}_{k}$ denote the language models specialized on each of the relevant datasets $k=1,\ldots,K$. The distributions $\widetilde{p}_{k},k=0,\ldots,K$ are imaginary and we do not intend to explicitly estimate them.
@@
The key assumption is that both Eq. equation 4 and equation 5 do not have context terms in the right-hand side quantities. See a weaker version of Assumption equation 5 in the appendix.
Goal: Our goal is identifying $\pi_{k},k=1,\ldots,K$. Which probability distance metrics help identify these contributions? We want to do this without explicitly estimating $\widetilde{p}_{k},k=1,\ldots,K$.
@@
For each of the $m$ queries, we perform $K+1$ prompts (or $2^{K}$ prompts) and write the results in a linear equation as follow:
$\bm{p}^{(j)}=\Pi\widetilde{\bm{p}}_{k}^{(j)},\quad\text{for}\quad j=1,\ldots,m.$
We observe the quantity on the left-hand side, but none of the quantities in the right-hand side. The matrix $\Pi$ has a specific structure. By defining the matrix $P=[\bm{p}^{(1)},\ldots,\bm{p}^{(m)}]$, we can write our problem in the following matrix form:
@@
where $P\in[0,1]^{(K+1)\times m}$, $\Pi\in[0,1]^{(K+1)\times(K+1)}$, and $\widetilde{P}\in[0,1]^{(K+1)\times m}$.
This is a matrix factorization problem with a special structure. We assume that $\widetilde{p}_{k}(y|q)$ can be obtained via some clever prompts. We can make assumptions about $\widetilde{p}_{k}(y|q)$ that allow recovery of $\bm{\pi}$ mixture parameters.
Remark 3: Instead of $K+1$ prompts, we can have up to $2^{K}$ prompts. However, for the prompts that use multiple datasets, we need to assume the form of the resulting distribution, similar to Eq. (5). An alternative is to impose priors on $\bm{\pi}$ and $\widetilde{P}$ to improve identifiability. We pursue the second approach in the next section.
@@
We can have multiple estimates for $\bm{\pi}$ from Eq. (6). We can resolve this issue be encouraging solutions that have lower variance. We achieve this by using two regularizers: an entropy regularizer for $\bm{\pi}$ to assume that the sources contribute equally and a regularizer that encourages $\widetilde{P}$ to be less informative.
@@
$\qquad-\lambda_{\pi}H(\bm{\pi})+\lambda_{\widetilde{P}}\|\widetilde{P}-1/2\|_{F}^{2},$
[…]
$\qquad 0\preceq\widetilde{P}\preceq 1,$
where $\|\cdot\|_{F}$ and $H(\cdot)$ denote the Frobenius norm and Shannon’s entropy. We use entropy regularization on $\bm{\pi}$ encouraging the null hypothesis of “equal contributions of all sources”. The Frobenius norm regularization implies that unless there is strong evidence, the outputs of the latent probabilities $\widetilde{P}$ should be $1/2$. Note the regularizers are vital for obtaining a non-trivial solution, and in absence of the them, there are many solutions for the problem.
The problem in Eq. equation 7 is biconvex; i.e., fixing either of $\bm{\pi}$ or $\widetilde{P}$, the problem is convex (Gorski et al., 2007). Thus, we solve it by the alternating least squares method. We describe the procedure in Algorithm 2 in Appendix C. We further assist the regularization terms by randomly initializing $\widetilde{P}$ to be around $1/2$ and $\bm{\pi}$ to be around $1/(K+1)$. We obtain the confidence intervals for both SCM and CMF by bootstrapping. In Appendix B we study the impact of regularization coefficients on the uncertainty of the estimation using synthetic datasets.
@@ 3.1 Prompt Engineering
For simplicity of evaluation and without loss of generality, we used BoolQ (Clark et al., 2019) Q&A dataset, where the answers are binary Yes/No. To instruct the LLMs to provide direct boolean responses, we used prompt engineering. Initially, we tested various prompts without explicitly instructing the model to answer with ”Yes” or ”No.” Diverse examples used in this process are provided in Appendix D.3. Through iterative testing, we found that responses improved when the model was explicitly instructed to provide a boolean answer. This led to our final prompt:
Prompt: ”Given the context below, answer the question that follows with only ’Yes’, ’No’, or ’I don’t know’ if the context is insufficient.
{question}? The answer to this question is”
While this final prompt worked well for GPT-4, Bloomz, and Mistral 7B, generating straightforward ”Yes,” ”No,” or ”I don’t know” responses, it was harder to instruct Phi-3-mini. Even with the final prompt, Phi-3-mini often generated more text than just a simple boolean response.
@@ 3.2 Using RAG
Given the limitations of LLM context windows, fitting entire datasets directly into the context is impractical. To address this, we utilized Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) to enhance the context by retrieving relevant documents from databases before generating responses. The process involves splitting the documents into semantically relevant chunks using the RecursiveCharacterTextSplitter from the Hugging Face Transformers library, computing embeddings for all chunks with a model like thenlper/gte-small, and storing these embeddings in a vector database using FAISS (Facebook AI Similarity Search). When a question is posed, it is embedded and a similarity search is performed against the vector database to find the closest matching documents. These retrieved documents are then provided as context to the LLMs along with the original question, allowing the LLMs to generate responses augmented with additional context.
We used a chunk size of 512 and a top-k value of 3, ensuring the context was trimmed to 2000 characters for conciseness.
4 Preliminary Experiments
@@
Step 1: Task Selection We use the BoolQ Q&A dataset, which consists of tuples in the form (question, relevant context, binary answer) for each question.
@@
Step 2: LLM Selection We examined four instruction-tuned LLMs: GPT-4, Bloomz, Mistral 7B, and Phi-3-mini. We report the accuracy of these LLMs on the Q&A task in Table 5. Given that the dataset is binary, we prompted the LLMs to answer ”Yes” or ”No” to each question, or to say ”I don’t know” if they could not provide a definite response (see Section D.1).
@@
Step 3: Dataset Collection We collected five datasets on different topics. The corpora were sourced from a subset of the Wikipedia Field of Science dataset available on Hugging Face, specifically the fields of Chemistry, Natural Science, History and Archaeology, Biology, and Law. Each of these datasets contains more than a million samples across five categories.
@@
Step 4: Evaluation Clearly, the context from BoolQ is more related to the questions. Thus, the successful methods should estimate higher weight for BoolQ, as the proxy for the relevant data used during training. For SCM, we report $\phi_{k}$ for 6 sources. Similarly, we report 7-dimensional $\bm{\pi}$ vector for CMF. The first dimension $\pi_{0}$ represents the contribution of all other data sources.
Results and Analysis

### v2

@@ Abstract
Fast Training Dataset Attribution via In-Context Learning
Milad Fotouhi, Mohammad Taha Bahadori, Oluwaseyi Feyisetan, Payman Arabshahi, David Heckerman Amazon and University of Washington {mfotouhi, bahadori}@uw.edu
@@ 1 Introduction
Training Data Attribution (TDA) refers to the task of quantifying the contributions of different data sources on the output of a model (Park et al., 2023; Nguyen et al., 2023). This task is essential for debugging the curating corpora processes for training and for improving the training of neural networks Xia et al. (2024); Qin et al. (2025). Understanding the contribution of data sources allows us to assess the monetary value of proprietary training data, which is crucial for fair compensation and data management (Ghorbani and Zou, 2019; Nohyun et al., 2022; Choe et al., 2024).
Existing TDA methods fall mainly into two categories: retraining-based methods and influence function-based methods, as detailed in recent surveys (Hammoudeh and Lowd, 2024; Worledge et al., 2024). Retraining approaches such as those of (Feldman and Zhang, 2020; Ghorbani and Zou, 2019) involve retraining the model without the target data source. However, this method is computationally expensive. The influence function approaches (Koh and Liang, 2017; Pruthi et al., 2020; Chen et al., 2021; Park et al., 2023), relax the need for full retraining by requiring only a few gradient calculations with respect to the data. Despite their efficiency, these methods rely on a linear approximation of the neural network around the target data point, which can be inaccurate. Critically, the influence function approaches compute the attribution score for a dataset as a linear function (usually an average or sum) of the attribution scores for each data point in the dataset (Hammoudeh and Lowd, 2024; Park et al., 2023). This approach fails to provide a holistic view of the contributions of an entire dataset to the model’s output. Additionally, both methods require access to the internals of LLMs, which is not feasible for some popular models. A related technique, Machine Unlearning (Ginart et al., 2019; Sekhari et al., 2021) is still expensive to obtain the contribution scores.
@@
In the experiments, we evaluated four instruction-tuned LLMs: Mistral 7B (Jiang et al., 2023), Bloomz (Le Scao et al., 2023), Microsoft/Phi-3-mini (Abdin et al., 2024) and GPT 4.0 (Achiam et al., 2023) on a set of binary Q&A datasets, BoolQ (Clark et al., 2019). In addition to the widely used BoolQ dataset, we create two new datasets: FakeQ, a synthetically modified version of BoolQ with altered queries and contexts, and a novel dataset constructed from Olympic 2024 Paris information, which serves as a realistic dataset that none of the LLMs have encountered during training.
Finally, to ensure the reliability of our proposed contribution estimation metrics, we fine-tune these LLMs in the Olympic 2024 dataset under varying conditions, such as different learning rates, and evaluate the consistency of the metrics. Once validated, the metrics are further used to assess and rank popular unlearning techniques. Furthermore, we employ the Trak evaluation framework (Park et al., 2023) as a baseline to benchmark the effectiveness and robustness of our methods.
@@ 2 Methodology
In our setting, we have tuples in the format of question, context, and outcome: $(q,c,y)$. Our LLM outputs $M(q|c)=p(y|q,c)$. When we do not use any context, we denote $c=\emptyset$. Our goal is to quantify the contributions of the training datasets $D_{1},\ldots,D_{n}$ in $p(y|q,c)$. We assume that we have a query set $Q=\{q_{1},\ldots,q_{m}\}$. For simplicity of notation, without loss of generality, we describe the methods for binary outcome $y\in\{0,1\}$.
We assume that we have $k,k=1,\ldots,K$, relevant datasets about a topic and we want to quantify their contributions in the generation of the outputs by our LLM.
@@ 2.1 The Non-parametric Approach: The Shapley Context Method (SCM)
The key idea of this approach is that if an LLM uses the information from the $k$th dataset, providing the $k$th dataset as a context will not change the output much. Conversely, if adding a dataset as context changes the output significantly, it is likely not used for the generation of the output. We define the following similarity scores:
@@
where $c_{k}$ is the context from the $k$th dataset.
Usually, the desired information can be found in multiple data sets (Ghorbani and Zou, 2019). To consider the impact of datasets in presence of other datasets, we define the following scores to be used in the Shapley formula (Shapley, 1953):
$s_{S}=\mathrm{sim}(y\,,y|c_{S}).$
@@
$\phi_{k}=\sum_{S\subseteq\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}}C_{S,K}(s_{S\cup\{D_{k}\}}-s_{S}),$ (2)
where $C_{S,K}=|S|!(K-|S|-1)!/K!$ are the normalization constants. This formula finds the residual increase in the similarity by including $D_{k}$, when we already have included another set $S\subseteq\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}$. The following Algorithem describes the details of our Shapley Context Method (SCM).
Input: An instruction-tuned LLM $M$ that outputs $y$ for each query $q$ and context $c$.
Input: A set of queries $Q=\{q_{1},\ldots,q_{m}\}$.
Input: A set of datasets that we need to compute their contributions $D_{1},\ldots,D_{K}$.
1 for $q\in Q$ do
2       Compute output without context: $y=M(q)$.; for $S\subset\{D_{1},\ldots,D_{K}\}\setminus\{D_{k}\}$ do
3             Use RAG to create contexts $c_{S}$ and $c_{S\cup\{D_{k}\}}$ from the datasets in $S$ and $D_{k}$.;
4
5            Compute the output with the context $y|c_{S}=M(q|c_{S})$.;
6
7            Compute the output with the context including $D_{k}$: $y|c_{S\cup\{D_{k}\}}=M(q|c_{S\cup\{D_{k}\}})$.;
8
9            Compute the similarities $s_{S}$ and $s_{S\cup\{k\}}$.;
10
11       end for
12      Use Eq. (2) to compute $\phi_{k}(q)$.
13 end for
Result: Return average $\phi_{k}$ over $m$ queries
Algorithm 1 Shapley Context Method (SCM)
@@ 2.2 The Semi-Parametric Approach: Context Mixture Factorization (CMF)
We propose a model to summarize the behavior of LLMs. Our model explicitly contains attribution scores and captures the entirety of the datasets used for its training. We use a mixture distribution approach, which defines:
@@
where $\widetilde{p}_{0}$ denotes a general-purpose language model and $\widetilde{p}_{k}$ denote the language models specialized on each of the relevant datasets $k=1,\ldots,K$. The distributions $\widetilde{p}_{k},k=0,\ldots,K$, are latent, and we do not intend to explicitly estimate them.
@@
The key assumption is that both Eq. (3) and (4) do not have context terms on the right-hand side quantities.
Goal: Our goal is to identify $\pi_{k},k=1,\ldots,K$. We want to do this without explicitly estimating $\widetilde{p}_{k},k=1,\ldots,K$.
@@
For each of the $m$ queries, we perform $K+1$ prompts (or the maximum $2^{K}$ prompts) and write the results in a linear equation as follows:
@@
where $P\in[0,1]^{(K+1)\times m}$, $\Pi\in[0,1]^{(K+1)\times(K+1)}$, and $\widetilde{P}\in[0,1]^{(K+1)\times m}$. We observe the quantity on the left-hand side, but none of the quantities in the right-hand side.
This is a matrix factorization problem with a special structure. We assume that $\widetilde{p}_{k}(y|q)$ can be obtained by some clever prompts. We can make assumptions about $\widetilde{p}_{k}(y|q)$ that allow recovery of the mixture parameters of $\bm{\pi}$.
Remark 3: Instead of $K+1$ prompts, we can have up to $2^{K}$ prompts. However, for the prompts that use multiple datasets, we need to assume the form of the resulting distribution, similar to Eq. (4). An alternative is to impose priors on $\bm{\pi}$ and $\widetilde{P}$ to improve identifiability. We will discuss the second approach in the next section.
@@
We can have multiple estimates for $\bm{\pi}$ from Eq. (5). We can resolve this issue by encouraging solutions that have lower variance. We achieve this by using two regularizers: an entropy regularizer for $\bm{\pi}$ to assume that the sources contribute equally and a regularizer that encourages $\widetilde{P}$ to be less informative.
@@
$-\lambda_{\pi}H(\bm{\pi})+\lambda_{\widetilde{P}}\|\widetilde{P}-1/2\|_{F}^{2},$
s.t. $\quad\bm{\pi}\succeq\bm{0},\quad\bm{1}^{\top}\bm{\pi}=1,\quad 0\preceq\widetilde{P}\preceq 1.$
Input: An instruction-tuned LLM $M$ that outputs $y$ for each query $q$ and context $c$.
Input: A set of queries $Q=\{q_{1},\ldots,q_{m}\}$.
Input: A set of datasets that we need to compute their contributions $D_{1},\ldots,D_{K}$.
1 for $q\in Q$ do
2       Compute output without context: $p(y|q,c_{0})=M(q)$.; for $k=1,\ldots,K$ do
3             Use RAG to create context $c_{k}$ from the dataset $D_{k}$.;
4
5            Compute the output with the context $p(y|q,c_{k})=M(q|c_{k})$.
6       end for
7      Use Eq. (2) to compute $\phi_{k}(q)$.
8 end for
9Build matrix $P$, where $P_{k,j}=p(y|q_{j},c_{k})$.;
10
11Solve Eq. (6) via alternating least squares and to compute $\widehat{\bm{\pi}}$.;
12
Result: Return the contribution vector $\widehat{\bm{\pi}}$
Algorithm 2 Context Mixture Factorization (CMF)
where $\|\cdot\|_{F}$ and $H(\cdot)$ denote the Frobenius norm and Shannon’s entropy. We use entropy regularization on $\bm{\pi}$ to encourage the null hypothesis of “equal contributions of all sources”. Regularization of the Frobenius norm implies that, unless there is strong evidence, the outputs of the latent probabilities $\widetilde{P}$ should be $1/2$. Note that regularizers are vital for obtaining a non-trivial solution and in the absence of them, there are many solutions to the problem.
The problem in Eq. (6) is biconvex; ie, fixing $\bm{\pi}$ or $\widetilde{P}$, the problem is convex (Gorski et al., 2007). Thus, we solve it by the alternating least-squares method. We describe the procedure in Algorithm 2 . We further assist in the regularization terms by randomly initializing $\widetilde{P}$ to be around $1/2$ and $\bm{\pi}$ to be around $1/(K+1)$. We can obtain the confidence intervals for both SCM and CMF by bootstrapping (Tibshirani and Efron, 1993).
@@ 3.1 Prompt Engineering
For simplicity of evaluation and without loss of generality, we used Q&A datasets, where the answers are binary Yes/No. To instruct the LLMs to provide direct boolean responses, we used prompt engineering. Initially, we tested various prompts without explicitly instructing the model to answer with "Yes" or "No." Diverse examples used in this process are provided in Appendix A. Through iterative testing, we found that the responses improved when the model was explicitly instructed to provide a Boolean answer. This led to our final prompt:
Prompt: "Given the context below, answer the question that follows with only ’Yes’, ’No’, or ’I don’t know’ if the context is insufficient.
{question}? The answer to this question is "
Although this final prompt worked well for GPT-4, Bloomz, and Mistral 7B, generating straightforward "Yes," "No," or "I don’t know" responses, it was harder to instruct Phi-3-mini. Even with the final prompt, Phi-3-mini often generated more text than just a simple boolean response.
@@ 3.2 Using RAG
Given the limitations of LLM context windows, fitting entire datasets directly into the context is impractical. To address this, we used Retrieval Augmented Generation (RAG) (Lewis et al., 2020) to enhance context by retrieving relevant documents from databases before generating responses. The process involves splitting the documents into semantically relevant chunks using the RecursiveCharacterTextSplitter from the HuggingFace Transformers library, computing embeddings for all chunks with a model like thenlper/gte-small, and storing these embeddings in a vector database using FAISS (Facebook AI Similarity Search) Johnson et al. (2019). When a question is posed, it is embedded, and a similarity search is performed against the vector database to find the closest matching documents. These retrieved documents are then provided as context for the LLMs along with the original question, allowing the LLMs to generate responses augmented with additional context. We used a chunk size of 512 and a top-k value of 3, ensuring the context was trimmed to 2000 characters for conciseness. We study the effectiveness of RAG in the Appendix C.
4 Experiments
@@
Step 1: Task Selection We use three datasets for our evaluation: (1) the BoolQ Q&A dataset (Clark et al., 2019), which consists of tuples in the form (question, relevant context, binary answer), representing a dataset that instruction-tuned LLMs are likely to have been trained on; (2) the FakeQ dataset, constructed by altering the queries and contexts in BoolQ to ensure the dataset remains unseen by the LLMs; and (3) The Olympic 2024 dataset, a newly created dataset based on the Paris 2024 Olympics (detailed in Appendix B), is designed to simulate real-world scenarios with Yes/No questions and relevant contexts. The latter two datasets allow us to evaluate the attribution metrics on datasets to which the LLMs have not been exposed during pretraining.
@@
Step 2: LLM Selection We examined four instruction-tuned LLMs: GPT-4, Bloomz, Mistral 7B, and Phi-3-mini. We report the accuracy of these LLMs on BoolQ in Table 9. Given that the dataset is binary, we prompted the LLMs to answer "Yes" or "No" to each question, or to say "I don’t know" if they could not provide a definite response (see Section 3.1).
@@
Step 3: Alternative Dataset Collection We collected five datasets on different topics. The corpora were sourced from a subset of the Wikipedia Field of Science dataset available on Hugging Face, specifically the fields of Chemistry, Natural Science, History and Archaeology, Biology, and Law. Each of these data sets contains more than a million samples in five categories.
@@
Step 4: Evaluation First, we evaluated the methods on the BoolQ dataset, which is closely related to the questions asked, providing a baseline for how well the methods estimate attribution for data sources that align closely with the LLMs’ pretraining. Successful methods should estimate a higher weight for BoolQ, as a proxy for the relevant data used during training.
To further test the robustness of the proposed methods, we created two new datasets: (1) FakeQ, derived by altering the queries and contexts in BoolQ to ensure it is entirely unseen by the LLMs, and (2) Olympic2024, a dataset based on the Paris 2024 Olympics, designed as a real-world dataset with binary Yes/No questions and relevant contexts that are guaranteed to be unseen by the LLMs (trained prior to 2024). These data sets allowed us to investigate how attribution metrics behave when the datasets have no prior exposure during LLM training.
To demonstrate the validity of the attribution metrics, we performed a series of experiments involving fine-tuning the LLMs on the Olympic2024 dataset with increasing learning rates or iterations. By fine-tuning the LLMs incrementally on a previously unseen dataset, we created a progression of models. For each fine-tuned model, we applied the contribution estimation methods and observed that the metrics increased monotonically, reflecting the LLMs’ growing reliance on the dataset.
Finally, we extend our evaluation to machine unlearning algorithms, leveraging the established metrics to rank well-known unlearning methods based on their ability to effectively reduce the contribution of a target dataset. By applying attribution metrics to LLMs subjected to unlearning processes, we assessed whether the influence of the targeted dataset was effectively diminished.
4.1 Results and Analysis
In general, SCM and CMF demonstrate their ability to effectively identify the most influential datasets, with CMF providing more robust attributions by accounting for noise and base contributions. We also performed an evaluation using Trak (Park et al., 2023) as a popular baseline. Trak provides a different perspective on attribution of data sets by scoring the impact of training data on model predictions. The Trak scores for Bloomz and Phi-3-mini in the BoolQ, FakeQ, and Olympic2024 datasets are shown in Tables 1, 2, and 3.
Table 1: Attribution values for the context on BoolQ
Algorithm Bloomz GPT-4 Mistral 7B Phi-3
SCM 0.48 0.59 0.57 0.50
CMF 0.63 0.62 0.61 0.59
Trak 0.61 0.58
Table 2: Attribution values for the context on FakeQ
Algorithm Bloomz GPT-4 Mistral 7B Phi-3
SCM 0.32 0.36 0.33 0.28
CMF 0.46 0.43 0.41 0.38
Trak 0.39 0.35
Table 3: Attribution values for the context on Olympic2024
Algorithm Bloomz GPT-4 Mistral 7B Phi-3
SCM 0.08 0.11 0.09 0.07
CMF 0.16 0.14 0.12 0.10
Trak 0.12 0.09